1. How did technology increase U.S. economic growth in the 1990s?
  2. Over the past twenty years an economy’s total output has grown from 1000 to 1300, its capital stock has risen from 2500 to 3250, and its labor force has increased from 500 to 575. All measurements are in real terms. Calculate the contributions to economic growth of growth in capital, labor, and productivity: a. assuming that aK = 0.3 and aN = 0.7. b. assuming that aK = 0.5 and aN = 0.5.
1
Expert's answer
2021-10-26T17:26:47-0400

1. The invention of technology in the US strengthened the economy and improved the standard of living thus increasing the economic growth in the country.

2. (a)

20 years ago. Today. Percent change

Y"=" 1000. 1300. 30%

K"=" 2500. 3250. 30%

N"=" 500. 575. 15%

"\Delta" A/A"=" a"_K" "\Delta" K/K"=" a"_N" "\Delta" N/N

"=" 30%"-" "(" 0.3"\times" 30%")-" 0.7"\times" 15%

"=" 30%"-" 9%"-" 10.5%

"=" 10.5%

Capital growth contributed is 9% "(" a"_K" "\Delta" K/K")" labour growth contributed 10.5% "("a"_N" "\Delta" N/N")", productivity growth was 10.5%

(b) "\Delta" A/A"=" 30%"-(" 0.5"\times" 30%")-" "(" 0.5"\times" 15%")"

"=" 30%"-" 15 %"-" 7.5%

"=" 7.5%

Capital growth contributed 15%"(" a"_K" "\Delta" K/K")" labour growth contributed 7.5%"(" a"_N" "\Delta" N/N")" productivity growth was 7.5%